# Key Words.: Gravitational lensing: strong – Methods: data analysis – Methods: observational – Galaxies: clusters: individual: Perseus1 Introduction

### v1

@@ 4 Visual inspection
The experts are anonymised at every stage of the analysis to avoid biasing the visual inspection results. We refer to the experts only by their expert ID, for example, “expert number 12”.
@@ 4.1 Team 1
The first stage of the visual inspection was completed by 20 out of the 22 experts that registered for Team 1. Figure 4 presents the number of lens candidates selected by each expert. We note that expert number 1 selected more sources for reinspection when compared to the mean and standard deviation of the other experts from Team 1, which selected sources for reinspection. Figure 5 presents the distribution of votes with and without expert number 1. We observe that the number of sources selected only by a single expert drops by half when excluding the outlier expert. Consequently, we removed the classifications from expert number 1. This allowed us to have a generous selection for the second stage, in which we included every source chosen by any of the remaining experts, while also keeping the sample reasonably small. The total number of sources selected for inspection in stage 2 was .
The second stage was completed by 17 out of the 22 experts. Figure 12 presents the number of votes per grade per expert. Similarly, Table 1 presents the average number of votes per grade for every stage that used the 1-by-1 sequential tool. Both Figs. 12 and 1 show that the experts of Team 1 were conservative when assigning the highest grade, A. This trend, however, was not the case for grades B and C, for which multiple experts assigned hundreds of votes in the sample. This indicates confusion among some experts for what constitutes a lens when the lensing features are not strikingly obvious. However, the experts with a large number of positive classifications were still a minority. Ultimately, the behaviour of all experts during the second stage was deemed acceptable and none of classifications were removed.
@@ 4.2 Team 2
The second stage of Team 2 was completed by 17 experts out of the 19 who registered. The grading behaviour of the experts is presented in Fig. 14. A few of the experts gave a large number of positive grades, one grading positively 673 out of the 691. Still, most of the experts reserved the grades A and B for the few best candidates, thus, we did not remove any experts. This is justified by the aggregation method introduced in Sect. 3.2, which relies on a majority vote to assign final grades.
@@ 4.3 Time cost of the inspections
It took the experts of Team 1 between and to inspect the whole parent sample using the mosaic tool, with the median time being . By contrast, it took the experts from Team 2 between to with a median time of to inspect one third of the parent sample, but using the 1-by-1 sequential tool. The median time per source was and for Teams 1 and 2, respectively.
For the second stage, Team 1 experts took between and . The median time was , equivalent to per source. Finally, Team 2 experts took between and , and the median was , which corresponds to per source. Overall, the time taken per source during the second stage was consistent for both teams. Given the filtering of the sample during the first stage, the time taken per source in the second stage corresponds more closely to the real time that a visual inspection would take on a preselected sample, e.g., the output of a CNN.
@@ 5.1 The visual inspection candidates
We compute three grades for each source: two individual grades from Team 1 and Team 2, plus a final joint grade. The final grade of each team is determined by applying the scheme presented in Sect. 3.2 to their respective set of votes. The joint final grade is determined the same way but using the votes of both teams together in the computation, with the exception of sources that were rejected by one of the teams during the first stage. In that case, we add two negative votes to the classifications of the other team and calculate the final joint classification using the 19 votes.
Overall, we obtain three grade A, 13 grade B, and 52 grade C lens candidates from the final joint grades. We present a summary of the number of candidates per team in Fig. 7. Furthermore, we show the grade A and B candidates in Fig. 8, along with their three final grades: the final grade from Team 1, the final grade from Team 2, and the final joint grade (in yellow, cyan, and red, respectively). The grade C candidates, along with the sources selected by either team but rejected from the joint sample, are presented in Appendix B in Figs. 15, 16, 17, 18 and 19. We note that Team 1 found almost double the amount of grade A and grade B candidates, but about the same number of grade C candidates as Team 2. Moreover, all the candidates selected by Team 1 but rejected from the final sample were also rejected by Team 2 during the first stage. By contrast, most of the sources selected by Team 2 but rejected from the final sample were also rejected by Team 1 during the second stage. This suggests that using the 1-by-1 sequential tool to filter out the obvious non-lenses, as Team 2 did during the first stage, is in fact a more aggressive filter than using the mosaic tool to preselect candidates.
@@ 5.2 Modelling of the lens candidates
We further assess the validity of the 16 grade A and B candidates by modelling the lens galaxy light and mass distribution, and surface brightness distribution of the lensed source. We do this using the pronto software (Vegetti & Koopmans 2009; Rybak et al. 2015a, b; Rizzo et al. 2018; Ritondale et al. 2019; Powell et al. 2021). We use a singular isothermal ellipsoid (SIE) density profile to describe the mass distribution of the galaxy, a composite of three Sérsic profiles for the lens galaxy light, and a pixellated, free-form reconstruction of the source surface brightness distribution. The parameters for the lens light and mass models are found through a non-linear optimisation using MultiNest (Feroz et al. 2009). For each set of light and mass parameters, the surface brightness distribution of the source is solved for linearly, up to some regularisation condition, which, in this case, penalises large gradients in the source. The strength of this regularisation is itself a non-linear parameter. The full model has 23 free parameters.
@@
We use a circular mask centred on the lensing galaxy and large enough to enclose what are assumed to be lensed images. We also use the positions of these assumed lensed images as an input to the optimisation scheme. After selecting 2, 3, or 4 positions in the image plane, the non-linear optimiser will only accept models where these image plane positions have an RMS separation in the source plane below some tolerance, in this case 1 arcsec. This removes the need to fine tune the initial modelling conditions, since only combinations of lens parameters which focus the source are accepted. The large tolerance of 1 arcsec also prevents the subjective choice of image positions from placing strict restrictions on the model.
After the optimisation, we check each model against three criteria to determine if its data are well described by a strong lensing model. These criteria are: (1) is there an SIE critical curve that can enclose or exclude the right number of bright components in the image plane; (2) is this critical curve, itself an isodensity contour, consistent with the lens galaxy light profile; and (3) is the reconstructed source surface brightness distribution consistent with a compact, focused object, inside a caustic.
@@
Five of the 16 grade A and B candidates are found to have valid lens models according to the criteria above. Six candidates cannot be determined as lenses with the available data, and a final five candidates do not have valid lens models. For the five valid candidates, we measure Einstein radii in the range –. The result for one valid candidate is shown, as an example, in Fig. 9, with the rest being shown in Appendix C.
Name | RA | Dec | Final joint grade | Modelling | Einstein Radius [arcsec] | ||
J031715+414057 | 49.315974 | 41.682676 | 20.810 | 0.004 | B | Valid | 0.70 |
J031741+412702 | 49.424405 | 41.450783 | 22.393 | 0.007 | B | Valid | 1.01 |
J031749+420011 | 49.456529 | 42.003291 | 19.791 | 0.003 | A | Valid | 0.92 |
J031953+415746 | 49.972097 | 41.962896 | 19.576 | 0.003 | A | Valid | 1.24 |
J031959+414229 | 49.999999 | 41.708073 | 22.467 | 0.007 | B | Valid | 0.68 |
J031615+414402 | 49.063959 | 41.733984 | 21.651 | 0.005 | B | Undetermined | |
J031640+414947 | 49.169013 | 41.829837 | 21.902 | 0.007 | B | Undetermined | |
J031647+414210 | 49.196262 | 41.702880 | 21.193 | 0.003 | A | Undetermined | |
J031648+415311 | 49.200924 | 41.886438 | 21.171 | 0.005 | B | Undetermined | |
J031834+413656 | 49.642982 | 41.615827 | 21.280 | 0.002 | B | Undetermined | |
J031838+413940 | 49.659347 | 41.661254 | 20.905 | 0.002 | B | Undetermined | |
J031646+414840 | 49.195286 | 41.811343 | 20.837 | 0.002 | B | Not valid | |
J031733+414944 | 49.388333 | 41.829079 | 20.666 | 0.002 | B | Not valid | |
J031925+420547 | 49.854480 | 42.096580 | 19.461 | 0.001 | B | Not valid | |
J031930+413711 | 49.877131 | 41.619889 | 22.147 | 0.004 | B | Not valid | |
J031941+414953 | 49.924888 | 41.831457 | 22.014 | 0.003 | B | Not valid |
@@ 6.1 The lens candidates in the literature
We crossmatch the sky coordinates of the 68 lens candidates against the Strong Lens Database (SLED, Vernardos et al., in prep.), a database of gravitational lenses, lens candidates, and known contaminants, encompassing more than entries. We found no matches between our sample and the database. Furthermore, the database indicates that there have been no lens searches in the area, since the closest match is more than away. This is expected given that the Perseus cluster is near the Galactic plane, and thus, is not generally targeted by large-scale optical galaxy surveys.
Nonetheless, Li et al. (2024) have performed a lens search in some of the ERO fields, including the Perseus cluster, in order to test the lens detection algorithm to be used by the Chinese Space Station Telescope (CSST) team. For that purpose, they used the high-resolution media images published by ESA.222Available here: https://www.esa.int/Science_Exploration/Space_Science/Euclid/Euclid_s_first_images_the_dazzling_edge_of_darkness They found four lens candidates in the Perseus cluster and reported the pixel coordinates in the TIFF image. Three of those candidates are included in our parent sample and were selected by Team 1 for reinspection in the second stage. However, all of them were rejected in the second stage. A smaller committee of experts inspected the fourth candidate and ultimately rejected it. Thus, we believe that none of the Li et al. (2024) candidates in the Perseus cluster are lenses. The lack of real lenses in their sample is likely a consequence of using the lower resolution TIFF images and the fact that their model was not trained to detect lenses in Euclid data. Therefore, the false positives in the Perseus cluster do not necessarily reflect on the potential performance of the algorithm in CSST data.
@@ 6.2 The expected prevalence of ELSE lenses
The median Einstein radius forecast in C15 is , whereas the smallest Einstein radius of the five lenses with valid models in Sect. 5.2 is . This tension is alleviated somewhat by accounting for the fact we only look at lenses with ¡ 23: by excluding the faintest lenses we prefer higher mass or lower redshift lenses, both of which result in larger Einstein radii. After applying this cut, the forecast median Einstein radius increases to . Figure 10 shows the expected number of lenses in of Euclid data, their Einstein radius distribution, and the Einstein radius distribution of successfully modelled ELSE candidates. The C15 forecasts assumed that the lens galaxy light would be subtracted, which is most helpful for finding small Einstein radius lenses: we have not done this. In future work, we will calibrate this selection function using simulations (Rojas et al., in prep.) but for now, we use a toy model to estimate it. We assume that the shape of the underlying Einstein radius population follows the Euclid forecasts of C15, but the visual inspection selection function is an additional step function where we discover every lens above a threshold Einstein radius (), and none below this threshold. Using Bayes’ Theorem, we infer given that we discovered five lenses and the lowest Einstein radius was in the range –.
We use a Monte Carlo simulation to draw many realisations of five Einstein radii, but vary the discovery threshold, assuming a uniform prior on the discovery threshold, . We find that the threshold is arcsec at 68% confidence. This corresponds to a forecast total population of lenses in the full Euclid data set that are discoverable with visual inspection without lens light subtraction and with lens ¡ 23. Moreover, scaling down the area to the forecast predicts approximately five lenses, coinciding with the five lens candidates with good models.
@@ 6.3 The prevalence of ELSE lenses
We extrapolate the five lens candidates with valid lens model found in the ERO Perseus field to the entire EWS. For this, we make three key assumptions:
@@
we have found all the discoverable lenses with within the ERO Perseus field;
@@
the average number density of sources with in the EWS is the same as in the ERO Perseus field: extended sources per ;333About of sources are cluster members and they occlude about of the sky. In terms of number of sources, the effect of the cluster balances out.
@@
We define a lens as discoverable if it can be identified via expert visual inspection and a valid lens model can be found for it. This assumption means that, in fact, we are mostly estimating the prevalence of larger Einstein radius lenses, which are easier to model and experts will identify more easily. However, small Einstein radius lenses can still be identified if the source is bright and the geometry of the lensed images is particularly evident; for example, the mock lens in the bottom right panel of Fig. 3, which has an Einstein radius of , but also a typical quad geometry, was still graded A by the experts.
@@
where and are the total numbers of lenses and sources in the EWS. Using the second assumption, we estimate the total number of extended sources with in the of the EWS to be . The number of trialled sources was , and the number of discoverable lenses was . We can then compute the posterior probability by Bayes’ theorem, i.e, . We assume that the prior probability of , , is such that the ratio is distributed uniformly in log-space between and unity. Overall, we expect ELSE-type lenses in the entire EWS. The posterior distribution is presented in Fig. 11, with the 68% confidence interval highlighted in light blue. Figure 11 also shows in orange the forecast of C15 as described in the previous section, which lies inside the confidence range of our estimate, showing a good agreement between the forecast and our estimate.

### v2

@@ 4 Visual inspection
The experts are anonymised at every stage of the analysis to avoid biasing the visual inspection results. We refer to the experts only by their expert ID, for example, ‘expert number 12’.
@@ 4.1 Team 1
The first stage of the visual inspection was completed by 20 out of the 22 experts that registered for Team 1. Figure 4 presents the number of lens candidates selected by each expert. We note that expert number 1 selected more sources for reinspection when compared to the mean and standard deviation of the other experts from Team 1, which selected sources for reinspection. Figure 5 presents the distribution of votes with and without expert number 1. We observe that the number of sources selected only by a single expert drops by half when excluding the outlier expert. Consequently, we removed the classifications from expert number 1. This provided us with a generous selection for the second stage, in which we included every source chosen by any of the remaining experts, while also keeping the sample reasonably small. The total number of sources selected for inspection in stage 2 was .
The second stage was completed by 17 out of the 22 experts. Figure 12 presents the number of votes per grade per expert. Similarly, Table 1 presents the average number of votes per grade for every stage that used the one-by-one sequential tool. Both Figs. 12 and 1 show that the experts of Team 1 were conservative when assigning the highest grade, A. This trend, however, was not the case for grades B and C, for which multiple experts assigned hundreds of votes in the sample. This indicates confusion among some experts for what constitutes a lens when the lensing features are not strikingly obvious. However, the experts with a large number of positive classifications were still a minority. Ultimately, the behaviour of all experts during the second stage was deemed acceptable and none of classifications were removed.
@@ 4.2 Team 2
The second stage of Team 2 was completed by 17 experts out of the 19 who registered. The grading behaviour of the experts is presented in Fig. 14. A few of the experts gave a large number of positive grades, one grading positively 673 out of the 691. Still, most of the experts reserved the grades A and B for the few best candidates, and thus, we did not remove any experts. This is justified by the aggregation method introduced in Sect. 3.2, which relies on a majority vote to assign final grades.
@@ 4.3 Time cost of the inspections
It took the experts of Team 1 between and to inspect the whole parent sample using the mosaic tool, with the median time being . By contrast, it took the experts from Team 2 between to with a median time of to inspect one third of the parent sample, but using the one-by-one sequential tool. The median time per source was and for Teams 1 and 2, respectively.
For the second stage, Team 1 experts took between and . The median time was , equivalent to per source. Finally, Team 2 experts took between and , and the median was , which corresponds to per source. Overall, the time taken per source during the second stage was consistent for both teams. Given the filtering of the sample during the first stage, the time taken per source in the second stage corresponds more closely to the real time that a visual inspection would take on a preselected sample, for example the output of a convolutional neural network.
@@ 5.1 The visual inspection candidates
We computed three grades for each source: two individual grades from Team 1 and Team 2 plus a final joint grade. The final grade of each team is determined by applying the scheme presented in Sect. 3.2 to their respective set of votes. The joint final grade is determined the same way but using the votes of both teams together in the computation, with the exception of sources that were rejected by one of the teams during the first stage. In that case, we added two negative votes to the classifications of the other team and calculated the final joint classification using the 19 votes.
Overall, we obtain three grade A, 13 grade B, and 52 grade C lens candidates from the final joint grades. We present a summary of the number of candidates per team in Fig. 7. Furthermore, we show the grade A and B candidates in Fig. 8, along with their three final grades: the final grade from Team 1, the final grade from Team 2, and the final joint grade (in yellow, cyan, and red, respectively). The grade C candidates, along with the sources selected by either team but rejected from the joint sample, are presented in a Zenodo appendix. We note that Team 1 found almost double the amount of grade A and grade B candidates, but about the same number of grade C candidates as Team 2. Moreover, all the candidates selected by Team 1 but rejected from the final sample were also rejected by Team 2 during the first stage. By contrast, most of the sources selected by Team 2 but rejected from the final sample were also rejected by Team 1 during the second stage. This suggests that using the one-by-one sequential tool to filter out the obvious non-lenses, as Team 2 did during the first stage, is in fact a more aggressive filter than using the mosaic tool to preselect candidates.
@@ 5.2 Modelling of the lens candidates
We further assessed the validity of the 16 grade A and B candidates by modelling the lens galaxy light and mass distribution, and surface brightness distribution of the lensed source. We did this using the pronto software (Vegetti & Koopmans, 2009; Rybak et al., 2015a, b; Rizzo et al., 2018; Ritondale et al., 2019; Powell et al., 2021). We used a singular isothermal ellipsoid (SIE) density profile to describe the mass distribution of the galaxy, a composite of three Sérsic profiles for the lens galaxy light, and a pixelated, free-form reconstruction of the source surface brightness distribution. The parameters for the lens light and mass models are found through a non-linear optimisation using MultiNest (Feroz et al., 2009). For each set of light and mass parameters, the surface brightness distribution of the source is solved for linearly, up to some regularisation condition, which, in this case, penalises large gradients in the source. The strength of this regularisation is itself a non-linear parameter. The full model has 23 free parameters.
@@
We used a circular mask centred on the lensing galaxy and large enough to enclose what are assumed to be lensed images. We also used the positions of these assumed lensed images as an input to the optimisation scheme. After selecting two, three, or four positions in the image plane, the non-linear optimiser will only accept models where these image plane positions have a root mean square separation in the source plane below some tolerance, in this case 1 arcsec. This removes the need to fine-tune the initial modelling conditions, since only combinations of lens parameters that focus the source are accepted. The large tolerance of 1 arcsec also prevents the subjective choice of image positions from placing strict restrictions on the model.
After the optimisation, we checked each model against three criteria to determine if its data are well described by a strong lensing model. These criteria are:
1.
Is there an SIE critical curve that can enclose or exclude the right number of bright components in the image plane?
2.
Is the centroid of this critical curve consistent with that of the light profile?
3.
Is the reconstructed source surface brightness distribution consistent with a compact, focused object, inside a caustic?
Given the small sample size, the evaluation of these criteria is done in a qualitative basis through the visual inspection of the models (see Appendix B). A systematic and quantitative assessment requires evaluating these criteria against realistic simulations where the ground-truth is known, which is beyond the scope of this work.
@@
Five of the 16 grade A and B candidates are found to have valid lens models according to the criteria above. Six candidates cannot be determined as lenses with the available data, and a final five candidates do not have valid lens models. For the five valid candidates, we measure Einstein radii in the range –. The result for one valid candidate is shown, as an example, in Fig. 9; the rest are shown in Appendix B. The results of the visual inspection and modelling are summarized in Table 2.
Name | RA | Dec | Grade | Modelling | [arcsec] | |
EUCL J031715.83+414057.6 | 49.315974 | 41.682676 | B | Valid | 0.70 | |
EUCL J031741.85+412702.8 | 49.424405 | 41.450783 | B | Valid | 1.01 | |
EUCL J031749.56+420011.8 | 49.456529 | 42.003291 | A | Valid | 0.92 | |
EUCL J031953.30+415746.4 | 49.972097 | 41.962896 | A | Valid | 1.24 | |
EUCL J031959.99+414229.0 | 49.999999 | 41.708073 | B | Valid | 0.68 | |
EUCL J031615.35+414402.3 | 49.063959 | 41.733984 | B | Undetermined | ||
EUCL J031640.56+414947.4 | 49.169013 | 41.829837 | B | Undetermined | ||
EUCL J031647.10+414210.3 | 49.196262 | 41.702880 | A | Undetermined | ||
EUCL J031648.22+415311.1 | 49.200924 | 41.886438 | B | Undetermined | ||
EUCL J031834.31+413656.9 | 49.642982 | 41.615827 | B | Undetermined | ||
EUCL J031838.24+413940.5 | 49.659347 | 41.661254 | B | Undetermined | ||
EUCL J031646.86+414840.8 | 49.195286 | 41.811343 | B | Not valid | ||
EUCL J031733.19+414944.6 | 49.388333 | 41.829079 | B | Not valid | ||
EUCL J031925.07+420547.6 | 49.854480 | 42.096580 | B | Not valid | ||
EUCL J031930.51+413711.6 | 49.877131 | 41.619889 | B | Not valid | ||
EUCL J031941.97+414953.2 | 49.924888 | 41.831457 | B | Not valid |
@@ 6.1 The lens candidates in the literature
We crossmatched the sky coordinates of the 68 lens candidates against the Strong Lens Database (SLED; Vernardos et al., in prep.), a database of gravitational lenses, lens candidates, and known contaminants, encompassing more than entries. We found no matches between our sample and the database. Furthermore, the database indicates that there have been no lens searches in the area, since the closest match is more than away. This is expected given that the Perseus cluster is near the Galactic plane, and thus, is not generally targeted by large-scale optical galaxy surveys.
Nonetheless, Li et al. (2024) have performed a lens search in some of the ERO fields, including the Perseus cluster, in order to test the lens detection algorithm to be used by the Chinese Space Station Telescope (CSST) team. For that purpose, they used the high-resolution media images published by ESA.222Available here: https://www.esa.int/Science_Exploration/Space_Science/Euclid/Euclid_s_first_images_the_dazzling_edge_of_darkness They found four lens candidates in the Perseus cluster and reported the pixel coordinates in the TIFF image. Three of those candidates are included in our parent sample and were selected by Team 1 for reinspection in the second stage. However, all of them were rejected in the second stage. A smaller committee of experts inspected the fourth candidate and ultimately rejected it. Thus, we believe that none of the Li et al. (2024) candidates in the Perseus cluster are lenses. The lack of real lenses in their sample is likely a consequence of using the lower-resolution TIFF images and the fact that their model was not trained to detect lenses in Euclid data. Therefore, the false positives in the Perseus cluster do not necessarily reflect on the potential performance of the algorithm in CSST data.
@@ 6.2 The expected prevalence of ELSE lenses
The median Einstein radius forecast in C15 is , whereas the smallest Einstein radius of the five lenses with valid models in Sect. 5.2 is . This tension is alleviated somewhat by accounting for the fact we only look at lenses with ¡ 23: by excluding the faintest lenses we prefer higher-mass or lower-redshift lenses, both of which result in larger Einstein radii. After applying this cut, the forecast median Einstein radius increases to . Figure 10 shows the expected number of lenses in of Euclid data, their Einstein radius distribution, and the Einstein radius distribution of successfully modelled ELSE candidates. The C15 forecasts assumed that the lens galaxy light would be subtracted, which is very helpful for finding small-Einstein-radius lenses: we have not done this. In future work, we will calibrate this selection function using simulations (Rojas et al., in prep.); in this study we used a toy model to estimate it. We assumed that the shape of the underlying Einstein radius population follows the Euclid forecasts of C15, and that the visual inspection introduced a selection function in the form of a step function, we detect every lens above a threshold Einstein radius () and none below this threshold. Using Bayes’ theorem, we infer given that we discovered five lenses and the lowest Einstein radius is in the range –.
We used a Monte Carlo simulation to draw many realisations of five Einstein radii but varied the discovery threshold, assuming a uniform prior on the discovery threshold, . We find that the threshold is arcsec at 68% confidence. This corresponds to a forecast total population of lenses in the full Euclid dataset that are discoverable with visual inspection without lens light subtraction and with lens ¡ 23. Moreover, scaling down the area to the forecast predicts approximately five lenses, coinciding with the five lens candidates with good models.
@@ 6.3 The prevalence of ELSE lenses
We extrapolated the five lens candidates with valid lens models found in the ERO Perseus field to the entire EWS. For this, we made three key assumptions:
@@
We have found all the discoverable lenses with within the ERO Perseus field.
@@
The average number density of sources with in the EWS is the same as in the ERO Perseus field: extended sources per .333About of sources are cluster members and they occlude about of the sky. In terms of number of sources, the effect of the cluster balances out.
@@
We defined a lens as discoverable if it can be identified via expert visual inspection and a valid lens model can be found for it. This assumption means that, in fact, we are mostly estimating the prevalence of larger Einstein radius lenses, which are easier to model and experts will identify more easily. However, small-Einstein-radius lenses can still be identified if the source is bright and the geometry of the lensed images is particularly evident; for example, the mock lens in the bottom right panel of Fig. 3, which has an Einstein radius of , but also a typical quad geometry, was still graded A by the experts. Furthermore, given the increased depth of the ERO Perseus field data compared to the EWS (0.75 magnitudes), we simulated cutouts for the 16 lens candidates with the typical S/N of the EWS. We observe that the lensing features are still visible and clearly identifiable in the shallower cutouts. Consequently, we assumed that they would had been identified in the shallower EWS.
@@
where and are the total numbers of lenses and sources in the EWS. Using the second assumption, we estimate the total number of extended sources with in the of the EWS to be . The number of trialled sources was , and the number of discoverable lenses was . We could then compute the posterior probability, , using Bayes’ theorem (i.e. ). We assumed that the prior probability of , , is such that the ratio is distributed uniformly in log-space between and unity. Overall, we expect ELSE-type lenses in the entire EWS. The posterior distribution is presented in Fig. 11, with the 68% confidence interval highlighted in light blue. Figure 11 also shows in orange the forecast of C15 as described in the previous section, which lies inside the confidence range of our estimate, showing a good agreement between the forecast and our estimate.